Find the unknown length using th hypotenuse theory.
a^2 + b^2 = c^2
b = 7, c = 12

Answer by richard1234(7193)

The Pythagorean "Theorem" is not a "theory," by the way.

a^2 + 7^2 = 12^2

a^2 = 12^2 - 7^2 = 144 - 49 = 95

a = sqrt(95)